Status Report

ISS On-Orbit Status 3 Jul 2002

By SpaceRef Editor
July 3, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below.

In the morning, CDR Valery Korzun and FE-2 Sergei Treschev completed
a scheduled inspection of the twelve porthole windows in the Service
Module (SM) and Docking Compartment (DC-1).  [Objective: to
evaluate window conditions in regard of chips, scratches, and
contamination on their surfaces.  The periodic inspection is
performed against a dark background (deep space), with the window
illuminated laterally by the Sun or directionally by a lateral light
source (portable lamp).  Any observed irregularities are
measured for extend and position, and MCC-M may then schedule still

FE-1 Peggy Whitson worked with the TOCA (total organic carbon
analyzer), processing the water samples from the hot water (SRV-K)
and supply water (SVO-ZV) ports collected yesterday.  [TOCA
is a water quality monitoring instrument for determining
concentrations of total organic carbon, total inorganic carbon, total
carbon, pH value and conductivity in water.  The lengthy
activity today was broken down in four segments, all using the same
procedure (detailed steps were uplinked).  The segmented
activities were separated by some unattended time (20 and 50 min),
and the TOCA hardware was stowed after the second 50-min period.]

Peggy Whitson also completed the scheduled maintenance of the
CSA-“>28compound specific analyzer-combustion products), which is
performed once every 30 days (to conserve battery life).
 [The CSA-CP is a passive cabin atmosphere monitor which
provides quick response capability during a combustion event (fire).
 Its collected data are stored on a logger.]

Korzun and Treschev worked another two hours unloading Progress
M-46 (8P) and stowing its cargo.  The activities are tracked
with the IMS inventory management system, which needs to be updated
on a daily basis.

The two cosmonauts completed their individual data sessions with the
NTXN psychological "Interactions" program, which keeps a
weekly record of their personal journal entries on mood and group
dynamics, including ground controllers, on an encrypted
questionnaire.  A parallel weekly activity is maintained by the
FCT (flight control team).

FE-1 Whitson conducted the regular weekly inspection of the RED
(resistive exercise device), followed today by the monthly procedure
of tightening the RED bolts.

Whitson also filled out the FFQ (food frequency questionnaire),
entering her food and fluid intake in a special nutrition tracking
program in the MEC (medical equipment computer).

In an attempt to recover the MCOR (medium-rate communications outage
recorder), which was powered off due to frequently reaching its upper
operational temperature limit, Peggy Whitson first checked out the
external fan for proper function, then removed and replaced the MCOR
docking station, followed by a checkout.  The MCOR is now
working properly.  The activity had been placed in the crew’s
"job jar" last night.

Another task list item, removal of launch restraint bolts of the CDRA
(carbon dioxide removal assembly), was also completed by the crew.

Daily routine servicing tasks performed today were SOSh life support
systems maintenance (Korzun), autonomous payload status check
(Whitson), and IMS "delta" file preparation (Treschev).

At 6:22 am EDT, the crew had a 10-min. amateur/ham radio exchange
with highschool students at Progymnasium Rosenfeld in Southern
Germany (Baden-WŸrttemberg), which is celebrating its 450th anniversary this year.  

New sessions of the Russian earth observation programs Uragan and
Diatomeya were added to Korzun’s and Treschev’s, to be performed

In the SM, TsVM (central computer system) lane 2 remains down, and
lane 1 is operating in stand-alone.  Restart of all three
mutually redundant lanes is still scheduled for 7/9 (Tuesday).
 Communication between TsVM and the U.S. C&C MDMs is

Today’s targets for the U.S. CEO (crew earth observation) program
were Taiwan Smog (the ISS passed directly over Taiwan, but
the site was be under the circulation of Typhoon Rammasun, forecast
to pass about 300 miles east of the island. Oblique views of this
120-kt storm may have been be possible well to the left of
, E. Mediterranean Dust and Smog (a cold front
pushing through the central Med was expected to continue to draw
Saharan dust northward ahead of it. On this descending pass, dust
should have been visible over the eastern Med from the Aegean Sea to
the Sinai)
, Rukwa Transform, Tanzania (this target was
likely to have some clouds at the time of the day of this pass, but
the low sun angle should have nicely enhanced the faults, folds, and
topographic structures of this relatively unsurveyed region.
 The ISS’ descending pass towards darkness cut the SW corner of
this target region, so crew was asked to try for off-nadir views to
the left of track)
, Angolan Biomass Burning (it is near
the height of fire season on southern Africa. On this descending
pass, the crew was to look obliquely to the left of track for sunlit
smoke plumes over the darkening surface below)
, Eastern United
(a significant smog event is now in progress over the
eastern U.S. Of interest on this pass: oblique and limb views of it
to the right of track)
, Eastern United States (high
pressure was causing a major buildup of smog over the eastern US. On
this pass, the crew was to look obliquely to the left of track to
document this event)
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 12:20 pm

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
 U.S. CDRA CO2 scrubber is Off (failed).  BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify

SM Working Compartment:  Pressure (mmHg) — 757,
temperature (deg C) — 28.1, ppO2 (mmHg) — 156.8, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 761,
temperature (deg C) — 20.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 20.7.
Node:  Pressure (mmHg) — 748.34, temperature (deg C) —
21.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 750.67 temperature (deg C)
— 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 750.67,
temperature (deg C) — 22.3; shell heater temp (deg C) — 22.6, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 22.5
PMA-2:  Shell heater temp (deg C) — 18.1

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) were replaced, but some more work
needs to be done.

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in
Directed position (non-solar tracking).
SM batteries: Battery #1 reported failed (6/21); all other batteries
(7) are in "Partial charge" mode.
FGB: Battery #6 is offline (failed); all other batteries (5) are in
“Partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 1 lane (of 3) operational.

Attitude Source:
3 CMGs online.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) was R&R’d today
and is operating nominally.

SSRMS/Canadarm2 at MBS PDGF (mobile base system/power &
data grapple fixture) and Lab PDGF, with Keep Alive power on both
strings (based on MBS).
MBS:  Operational on Redundant string, keep-alive power on Prime
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of this morning, 7:53 am EDT [=
Mean altitude — 391.8 km
Apogee — 396.9 km
Perigee — 386.7 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007498
Orbits per 24-hr. day — 15.58
Altitude decrease — 125 m (mean) in last 24 hours
Solar Beta Angle —  -17.5 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20657
Current Flight Attitude — LVLH (local vertical/local horizontal =
“earth-fixed”: z-axis in local vertical, x-axis in velocity vector
[yaw: -10 deg, pitch: 7.8 deg, roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

Everyone — have a happy 4thand a safe &
joyful holiday weekend!

SpaceRef staff editor.